OpenStudy (anonymous):

A = 21,500(1 + 0.06/2)^(2*6) WHAT IS THE ANSWER THANKS

OpenStudy (anonymous):

10769539847.86

OpenStudy (anonymous):

follow BEDMAS in order B= brackets E = exponent D=devision M=multiplication A=addition S=subtraction

OpenStudy (anonymous):

ZERONET I AM CONFUSED NOT SURE WHAT YOU ARE DOING BUT IT IS NOT HELPING ME WITH THE ANSWER

OpenStudy (anonymous):

a= 21500 (1+0.03)^(12) a= 21500 (1.03)^12 a= 21500 (1.42576...) a=

OpenStudy (anonymous):

Superstat I am not sure what the answer is I am not good at Math at all what is the correct answer?

OpenStudy (anonymous):

yeah i punched it in calc wrong. but regardless... follow BEDMAS rule

OpenStudy (anonymous):

just like tutordiffeq has written step by step

OpenStudy (anonymous):

\[A=30653.9\]